A population consists of the numbers 1, 2, 3, and 4 with the sample size of 2. Compute the variance of the sampling distribution of means.




1
Expert's answer
2022-04-15T04:58:06-0400

Population mean:

"\mu=\cfrac{1+2+3+4} {4} =2.5."

Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 1-2.5, 2-2.5, 3-2.5, 4-2.5
\end{Bmatrix}="

"=\begin{Bmatrix}
-1.5, -0.5, 0.5, 1.5
\end{Bmatrix},"

"\sigma^2=(-1.5)^2\cdot \cfrac{1}{4}+(-0.5)^2\cdot \cfrac{1}{4}+\\+0.5^2\cdot \cfrac{1}{4}+1.5^2\cdot \cfrac{1}{4}=1.25."


Variance of the sampling distribution of means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{1.25}{2}=0.625."